Problem 2:

The following events apply to Sam’s Seafood Restaurant for the year ended December 31, 2020, its first year of operations:

  1. The company acquired $50,000 cash by issuing common stock.
  2. Purchased a new cook top that cost $35,000 cash.
  3. Earned $36,000 in cash revenue.
  4. Paid $12,000 cash for salaries expense.
  5. Recorded depreciation expense on the cook top for 2020 using straight-line depreciation. The cooktop was purchased on January 1, 2020, the expected useful life of the cook top is four years, and the estimated salvage value is $3,000.

Required: Answer the following questions.

  1. What is the net income for 2020?             

  1. What amount of depreciation expense would Sam’s report on the 2021 income statement?

  1. What amount of accumulated depreciation would Sam’s report on the December 31, 2021, balance sheet?

  1. Would the cash flow from operating activites be affected by depreciation in 2021?
  1. If Sam’s Seafood Restaurant decided to sell the new cooktop in 2022 for $10,000, would the company realize a gain or loss? How much?
Answer #1

a) Net income = Revenue-Expense

Depreciation expense = (35000-3000/4) = 8000

Net income = 36000-12000-8000 = $16000

b) Depreciation expense = (35000-3000/4) = $8000

c) Accumulated depreciation = 8000*2 = $16000

d) Depreciation expense is non Cash expense so Cash flow from operating activities would not be affected by Depreciation

e) Gain (loss) = Sale value-Book value = 10000-(35000-8000*3) = -1000